Find the derivative of the function. 4x-9 y= 7x+4 The derivative i
Answer #1

Given function ,  = 4 -9 x+4

Differentiating both side with respect to x we get ,

(7.+ 4)是(4.7 -9)-(4-9)(7.2 + 4) (7z + 412

=y= 4(7.0 + 4) – 7(4.– 9) (7x + 4)2

y=- 28.0 + 16 - 28.6 +63 (7.x +4)2

79 y = 177 + 4)2

Hence required ,  79 y = 17.+ 4) .

Answer :  79 y = 172 + 4)

Here we have used the formula ,  () _ 2 (In) - 1 (12)

.

.

.

.

Please comment if you have any doubt .

RATE THE ANSWER ACCORDINGLY .